A simply supported one-way slab of 12 ft span and 8-inch thick is given. The rebar used in the slab is #8 at 6 inches on center. Determine how much live load (psf) may the slab safely carry. (f'c=3,000 psi, fy=40,000 psi).
Answer #1

Step 1:

L=12 ft

H=8 in

fc=3 ksi

Fy=40 ksi

Step 2:

As=#8 bars 6 in on center

As=1.57 in2/ft

As=bd

Consider d=8-2=6 in

1.57=x12x6

=0.0218

Step 3:

Mu/0.9x12x6x6=759.8 (from appendix A-TABLE-A-13)

Mu=295410=24.61 ft-kips

Mu=WuL2/8

24.61=Wux12x12/8

Wu=1.36 kips/ft is the maximum load possible on the slab